# arXiv PreprintSubmitted to arXiv on February 6, 2025

### v2

@@ 7 Conclusion and Implications
Understanding the EnKF as an empirical Matheron update opens up several possibilities, as both Matheron updates and Ensemble Kalman methods have been studied extensively in their respective domains. Both have witnessed many developments in the Bayesian Machine learning setting (Alzraiee et al., 2022; Chada and Tong, 2022; Chen et al., 2021, 2023; Dunbar et al., 2022; Guth et al., 2020; Huang et al., 2022; Kovachki and Stuart, 2019; MacKinlay et al., 2025; Oliver, 2022; Schillings and Stuart, 2017; Schneider et al., 2022; Spantini et al., 2022). If we understand the pathwise update of the Matheron in the setting of the empirical measures used in Ensemble Kalman filters, this suggests the potential to import developments from each of these to the other.
The literature of the EnKF is vast, and we have only scratched the surface of the many variants and extensions that have been proposed. Many of the ideas developed for the EnKF are thus potentially applicable to Gaussian Process regression via the Matheron update. For example, it provides a clear probabilistic interpretation of the ensemble update that may inspire improved regularization and localization strategies. For one example, the Local Ensemble Transform Kalman Filter (Bocquet et al., Tue Sep 22 00:00:00 UTC 2020) provides state-of-the-art performance in high-dimensional filtering and system identification settings, and could be adapted to the Matheron update.
@@ Appendix A Measures as densities
[…]

### v3

@@ 9 Conclusion and Implications
The affine residual identity is a Rosetta stone linking kriging, optimal-interpolation, ensemble-Kalman, and path-wise GP sampling. This synthesis upgrades each field: geoscientists inherit differentiable, variational, and sparse-GP tools; machine-learning researchers gain decades of localisation, inflation, and convergence theory; and inverse-problem analysts get a unifying algebra that bridges finite-ensemble and infinite-dimensional limits.
The literature of the EnKF is vast, and we have only scratched the surface of the many variants and extensions that have been proposed. Many of the ideas developed for the EnKF are thus potentially applicable to Gaussian Process regression via the Matheron update. For example, it provides a clear probabilistic interpretation of the ensemble update that may inspire improved regularization and localization strategies. For one example, the Local Ensemble Transform Kalman Filter (Bocquet et al., 2020) provides state-of-the-art performance in high-dimensional filtering and system identification settings, and could be adapted to the Matheron update.
@@
[…]
The author thanks CSIRO’s Machine Learning and Artifical Intelligence Future Science Platform for supporting this research.
The original of this contentwas drafted by human hands, with the exception of the experiments section and supporting code which was generated by LLMs.
@@ Appendix A Measures as densities
for any test function $f$. Here, the angle brackets $\langle\cdot,\cdot\rangle$ denote the action of the distribution on $f$. This sifting property ensures that when integrating a test function against the empirical density, we recover the sample average: